A disk takes 1 ms per track seek time , and the is head is currently on track 70. Assume it takes 5 ms to service a request on the head is at right track. The following set of requests are pending , given as (ID , track) in order of arrival. How much time does it take to service all the requests if the elevator algorithm is used and the current direction is down?

Requests ( A, 80) , (B,75) , (C,82) , (D,90) , (E,45)

Total time?
Answer #1

Given,

Track seek time = 1ms

Service request time = 5 ms

Elevator algorithm is referred to the SCAN algorithm.

Number of movements = | 70 - 45 | +  | 45 - 0 | +  | 0 - 75 | +  | 75 - 80 | +  | 80 - 82 | +  | 82 - 90 |

= 35 + 45 + 75 + 5 + 2 + 8

= 70 + 80 + 10

= 160

Time taken for seek = 160 * 1 ms = 160 ms

Time taken for servicing request = Number of tracks * Time for each track = 5*6 = 30

Total time = 160 + 30 = 190 ms.